OpenStudy (anonymous):

Use the slope and y-intercept method to graph the following equations. a.)3x-4y-1=27 b.)6x=2y+12

OpenStudy (anonymous):

plug in x=0 for y intercepts. plug y=0 for x intersepts then dray lines through the x and y intercepts...

OpenStudy (anonymous):

a) 3(0)-4y-1 = 27.......-4y = 28.......y=-7 so you have the point (0,-7)

OpenStudy (anonymous):

3x-4(0) = 28......x=28/3...... and now you have the point (28/3,0) so you can plot the points and draw a line through them because they are linear equations... do the same for b.